Axolotl:

What is the solution to the following system of equations? x − 3y = 5 2x + y = 10 (5, 0) (0, 5) (7, 0) (0, 7)

Vocaloid:

There’s a couple of ways to approach this, but the one I see off the bat is: you can multiply the second equation by 3 to get 3y From there, you can add the two equations. The -3y and the + 3y will then cancel out and you can solve for x

lifewmylilweeb:

what do you think it is?

Axolotl:

this one I am having trouble figuring out

Vocaloid:

Try what I suggested: can you multiply the whole second equation by 3 and lmk what you get?

Vocaloid:

The second equation is: 2x + y = 10 Notice what happens when we multiply everything on both sides by 3 3*2*x + 3*y = 3*10 Equals 6x + 3y = 30 Now our system is: x − 3y = 5 6x + 3y = 30 Can you try adding the two equations together from here?

Axolotl:

I still don't understand :/

Vocaloid:

The equations: x − 3y = 5 6x + 3y = 30 Add the x’s: x + 6x = 7x Add the y’s: -3y + 3y = 0 Add the constants 5+ 30 = 35 So the system is now 7x = 35 X = 5